Sanjeev scored 90% marks in mathematics, 85% Science, 80% Social Studies, and 75% Language.
If the maximum marks in mathematics is 200, science 150, Social Studies 150, and language 100, find the everybody at every great percentage if he wants to improve his aggregate percentage by 20% through language alone find the score required language

Answer by Theo(13342)


perfect score in math is 200.
perfect score in science is 150
perfect score in social studies is 150
perfect score in language is 100.

90% in math = .9 * 200 = 180.
85% in science = .85 * 150 = 127.5
80% in social studies = .8 * 150 = 120
75% in language = .75 * 100 = 75

total score is 180 + 127.5 + 120 + 75 = 502.5 out of a possible 200 + 150 + 150 + 100 = 600

502.5 / 600 = .8375 = 83.75%

to improve this by 20%, he will need to get an overall aggregate percentage of 1.2 * 83.75% = 100.5%.

he cannot improve his overall percentage by 20% because that would give him a score that is over 100%.

if he improved his language scores by 20%, then he would get 90 right instead of 75.

that would be an improvement of 15 points.

his overall score would then be 502.5 + 15 = 517.5.

his overall percentage would then be 517.5 / 600 = .8625 which is equal to 86.25%.

the increase in his overall percentage would be 86.25% minus 83.75% = 2.5%.

this represents a percent increase of 2.5% / 83.75% = .0298507463 = 2.98507463% over his original overall percentage.

his original percent was 502.5/600 = .8375 = 83.75%

his new percent would be 517.5/600 = .8625 = 86.25%.